OpenStudy (coolness):

NEW HELP MEDAL !!!!! EASY!!!!!!! HELP HELP HELP Which equations are related equations to x+14=12x+14=12? Select each correct answer. A. x=1214x=1214 B. x+14−14=12−14x+14−14=12−14 C. x+12=14x+12=14 D. x+14−14=12+14x+14−14=12+14 E. x=1134x=1134 F. x=12−14

OpenStudy (coolness):

@Ben00 @mathmale

OpenStudy (coolness):

@phi @paki @PandaNinja101 @thomaster

OpenStudy (coolness):

@FaiqRaees

OpenStudy (sapphiremoon):

What do you mean by "related"? Do you mean which equations are the same?

OpenStudy (coolness):

yes @SapphireMoon so the answer to x+1/4 = 12 is 3

OpenStudy (coolness):

Which equations are related equations to x+1/4=12 Select each correct answer. x=12 1/4x x+1/4−1/4=12−1/4 x+12=1/4 x+1/4−1/4=12+1/4 x=11 3/4 x=12−1/4

OpenStudy (coolness):

sorry this is what it should look like

OpenStudy (ben00):

Maybe x=12 1/4x???

OpenStudy (coolness):

but how @Ben00 if the answer to the equation is 3

OpenStudy (sapphiremoon):

@coolness I still can't make sense of these equations. For the first one, do you mean: \[x = 12 \times \frac{1}{4}x\]

OpenStudy (coolness):

sorry that x shouldnt be at the end of A @SapphireMoon

OpenStudy (coolness):

idk thanks anyways

OpenStudy (sapphiremoon):

Oh, does that mean 12 1/4 as in 12.25?